### v1

Abstract
Replacing non-polynomial functions (e.g., non-linear activation functions such as ReLU) in a neural network with their polynomial approximations is a standard practice in privacy-preserving machine learning. The resulting neural network, called polynomial approximation of neural network (PANN) in this paper, is compatible with advanced cryptosystems to enable privacy-preserving model inference. Using “highly precise” approximation, state-of-the-art PANN offers similar inference accuracy as the underlying backbone model. However, little is known about the effect of approximation, and existing literature often determined the required approximation precision empirically.
In this paper, we initiate the investigation of PANN as a standalone object. Specifically, our contribution is two-fold. Firstly, we provide an explanation on the effect of approximate error in PANN. In particular, we discovered that (1) PANN is susceptible to some type of perturbations; and (2) weight regularisation significantly reduces PANN’s accuracy. We support our explanation with experiments. Secondly, based on the insights from our investigations, we propose solutions to increase inference accuracy for PANN. Experiments showed that combination of our solutions is very effective: at the same precision, our PANN is 10% to 50% more accurate than state-of-the-arts; and at the same accuracy, our PANN only requires a precision of $2^{-9}$ while state-of-the-art solution requires a precision of $2^{-12}$ using the ResNet-20 model on CIFAR-10 dataset.
1 Introduction
Machine learning (ML) is revolutionising different industries in recent years. Despite its widespread adoption, a key challenge in deploying ML is ensuring data privacy. One promising direction is to employ advance cryptosystems such as multiparty computation (MPC) and homomorphic encryption (HE) during model training and model inference, resulting in privacy-preserving training and privacy-preserving inference. In this paper, we use the term Privacy-preserving machine learning (PPML) to cover both cases. While in principle these advance cryptosystems can be used to protect sensitive data used in the computation of any functionality, in practice they are designed to support basic arithmetics operations (i.e., addition and multiplication) or basic Boolean operations (i.e., AND and OR) only. This limitation makes it expensive to compute non-polynomial functions such as ReLU, sigmoid, and maxpool, functions commonly found in ML, and in particular, neural networks. Existing PPML address this limitations by replacing these functions with their polynomial approximations. In this paper, we refer to this type of neural network in which non-polynomial functions have been replaced with polynomials as polynomial approximation of neural network (PANN). Similar to existing MPC and HE based PPML, this paper focuses on privacy-preserving inference since privacy-preserving training using HE and MPC are still considered too inefficient in practice.
Figure 1: Time cost for our models and the state-of-the-art to achieve the previous best accuracy on PANN (CIFAR-10).
Naturally, the (implicit) use of PANN in privacy-preserving inference involves two design considerations, namely, the precision of the polynomial approximation and the accuracy of the PANN (compared with the underlying backbone model). Intuitively, the accuracy of inference cannot be guaranteed if the approximation introduces a large error. Early schemes often adopt simple polynomials in approximation. For example, CryptoNets and HCNN utilized the square function to replace activation functions Gilad-Bachrach et al. (2016); Badawi et al. (2020). CryptoDL used degree 2 and degree 3 polynomials to approximate the ReLU function Hesamifard et al. (2017), while Faster Cryptonets leveraged minimax approximation with degree two Chou et al. (2018). nGraph-HE and Delphi adopt quadratic approximations to approximate activation functions Boemer et al. (2019); Mishra et al. (2020). These methods were proposed to handle only simple model, as non-trivial errors generated by approximation result render the inference results useless in more complex neural networks. On the other hand, the use of high precision approximation guarantees inference accuracy at the cost of higher computational overhead. A common method for precise approximation is Minimax approximation Lee et al. (2022a; 2021b), which combines many small-degree polynomials to approximate non-polynomial functions. Although precise approximation can attain the same accuracy as the backbone model, it leads to much higher costs. Table 1 gives the time cost of the backbone model, their polynomial approximation at various precision, and the corresponding accuracy using state-of-the-art techniques. It can be seen that the PANN on ResNet-20 with precision $2^{-12}$ takes 71.2s to perform inference on the entire CIFAR-10 test set, whereas precision $2^{-8}$ only needs 28.2s. However, the inference accuracy is also significantly lower.
We make two additional remarks here. Firstly, while the overhead of using PANN is high, representing the non-polynomial functions using basic arithmetic or boolean operations is even more expensive. Secondly, we consider PANN as a standalone object, and compare their performance directly. When they are used in combination with advance cryptosystems, the difference will be further enlarged [^1] For instance, in MPC, the function to be computed is first converted into a boolean or arithmetic circuit. If the circuit consists of $n$ gate, the resulting cost of the MPC is $O(n)$ but not necessarily linear in $n$..
Some efforts have been made to increase PANNs’ accuracy while reducing overheads by customizing the model for PANN. For example, CryptoNets and Faster Cryptonets directly train models with low-degree polynomials approximation Gilad-Bachrach et al. (2016); Chou et al. (2018). AutoFHE fine-tunes the model on precise PANN to achieve better performance Ao & Boddeti (2023). However, there are several subtleties involved. Firstly, the polynomials’ derivative is unbounded, making it hard to run gradients descent. Secondly, these models are not generic. Different approximation methods or even different precision for the same method can use different polynomials. This implies that models must be trained separately for each method and precision. Thirdly, in many cases pre-trained model are available, and thus training (or fine-tuning) over PANN can result in much higher training cost and may not be desirable for real-world deployment.
Table 1: Time cost of backbone models (bb) and PANN with different precisions. Numbers in brackets indicate inference accuracy.
$2^{-8}$ $2^{-9}$ $2^{-10}$ $2^{-11}$ $2^{-12}$ bb
ResNet-20 28.2s 33.6s 42.3s 62.5s 71.2s 2.6s
(65.70) (89.69) (91.24) (91.42) (91.52) (91.56)
Shufflenetv2 38.1s 44.8s 50.5s 79.8s 89.7s 3.3s
(11.94) (32.08) (84.96) (87.24) (88.13) (88.60)
DLA-34 128.8s 156.1s 163.5s 251.8s 387.1s 5.3s
(12.78) (45.67) (88.53) (92.73) (94.45) (95.10)
Mobilenetv2 221.6s 282.7s 295.5s 479.1s 542.9s 3.9s
(10.78) (15.41) (82.85) (89.54) (90.90) (91.45)
To sum up, past researches appear to indicate that a trade-off seems inevitable: Low precision yields high efficiency but low accuracy. High precision yields high accuracy but unaffordable computation costs. The effect of approximation is relatively less understood. This serves as the motivation of this paper. Looking ahead, to better address this problem, as a first step, we separate PANN from the underlying cryptosystem, and investigate how approximation error affects inference accuracy in PANN. Based on this new understanding, we design a more general and efficient methodology to train backbone models so that the corresponding PANN gives reasonable inference accuracy even with approximation at low precision. Furthermore, we highlight that backbone models are still trained with non-polynomial components before they are used to produce PANN. The backbone models trained based on our method do not degrade its inference accuracy.
1.1 Overview of Our Results
We initiate the investigation on how to reduce the impact on inference accuracy due to approximation errors in PANN. For the ease of writing, we say a neural network $\mathbb{F}$ is sturdy if the PANN of $\mathbb{F}$ has good resistance against approximation errors. In other words, we can use less precise approximation for a sturdy neural network to maintain the inference accuracy. Our investigations lead to the following two interesting findings: (1) The concept of ”sturdiness” is similar to the notion of adversarial robustness (introduced by Projected Gradient Descent Attack (PGD) Kurakin et al. (2016)). Informally speaking, adversarial robustness aims to ensure the neural network is robust to perturbations of the inputs by an adversarial intent on misleading the classifier during the inference stage. We identified the similarity (and difference) between “sturdiness” and “robustness”. In particular, adversarial robustness is related to input perturbation and “sturdiness” is related to perturbation introduced by approximation errors. There are two differences: in adversarial robustness, perturbations are introduced intentionally at input, while for “stuidiness”, perturbations are introduced at every layer with approximation functions by approximation errors, which we referred to as intra-model perturbations. (2) Weight regularization commonly used during the training phase is extremely harmful to model “sturdiness”. Our empirical study shows the following counter-intuitive example: there exists a neural network $\mathbb{F}_{1}$ whose inference accuracy is higher than that of $\mathbb{F}_{2}$, yet the PANN of $\mathbb{F}_{2}$ performs much better than $\mathbb{F}_{1}$ at the same approximation precision. The only difference between $\mathbb{F}_{1}$ and $\mathbb{F}_{2}$ is the use of weight regularization during training of $\mathbb{F}_{1}$.
Based on our findings, we present two solutions to improve “sturdiness” of neural networks. Our first solution is based on the insight that “sturdiness” and robustness are similar in nature. In more detail, our first solution makes use of the technique known as adversarial training (AT), which increases robustness by introducing perturbation during training. However, naively applying perturbations at every layer requiring approximation does not work since it significantly damage the quality of the model. We address this issue by carefully crafting the kind of perturbation to be introduced during training. More concretely, due to the characteristics of the ReLU function, introducing perturbations on the negative part of the inputs during training will barely damage the accuracy of the backbone model. The reason is that information in these negative inputs should be filtered by ReLU function thus they are unimportant to the inference results. Yet, in PANN, these perturbations mimic the “intra-model” perturbations due to the approximation error. Since the perturbations are introduced at the negative part during training and usually corresponds to background information of inputs, we called them as perturbations on irrelevant information in the input background. And we called our method AT-like since it is inspired by adversarial training. Our second solution is based on the findings that weight regularization is harmful to “sturdiness”. Intuitively, we can improve “sturdiness” by removing weight regularization during training. However, weight regularization is crucial in providing generalization and training without it will results in a model with low accuracy. Our second solution is considered a trade-off: we use minimal weight regularization and combine it with Mixup to maintain an acceptable accuracy in the backbone model yet greatly increases its “sturdiness”.
Our two solutions are orthogonal and can be used together during training. We conduct extensive experiments to illustrate the effectiveness of our solutions. In particular, we showed that the resulting PANNs can achieve similar accuracy as state-of-the-art PANNs at a much lower approximation precision.
Our contributions are summarised below:
  • We initiate the study on PANNs and identified similarities and differences between the notion of adversarial robustness and “sturdiness”.
  • We observe that weight regularization can significantly reduce the “sturdiness” of neural networks, thus reducing the corresponding PANN’s accuracy.
  • We propose two solutions to enhance “sturdiness”, namely, an AT-like training method and reduce the use of weight regularization. Combined use of our solutions can significantly enhance “sturdiness”. Our solution is proven effective on various models and datasets. For instance, we achieve the state-of-the-art accuracy on CIFAR-10 with only 40% to 60% time cost (as shown in Figure 1).
1.2 Outline
The rest of the paper is organized as follows. Section 2 describes the preliminary of our study. In Section 3, we conduct an in-depth analysis on approximation errors in PANN. Section 4 provides solutions to improve low-precision PANN’s accuracy. Section 5 gives the experiment results. Finally, in Section 6, we conclude this paper and discuss potential avenues for future research.
2 Preliminary
2.1 Notation
We denote a neural network as $\mathbb{F}$, its approximation as $\widetilde{\mathbb{F}}$, benign inputs and label as ($x$, $y$). Perturbations for adversarial samples are represented by $\delta$. A small value is represented by $\epsilon$. The step length for PGD attacks is represented by $\alpha$. The precision for PANN is $2^{-\beta}$ determined by $\beta$. The approximation of $\operatorname{sgn}$ and $\operatorname{ReLU}$ are denoted by $\operatorname{Appsgn}$ and $\operatorname{AppReLU}$ respectively.
2.2 Polynomial Approximation of Neural Networks
Common approach for approximation in PANNs include simple polynomials, Taylor polynomials, Minimax approximation, etc. Among them, Minimax approximation is the most common because it can approximate for an interval and is computationally efficient. Therefore, our focus is primarily on studying PANN with Minimax approximation.
This paper refers to $p(z)$ as a polynomial approximation function. An $\beta$-close polynomial approximation of a function $f(z)$ in interval $[a,b]$ satisfies:
$|p(z)-f(z)|\leq 2^{-\beta},\quad x\in[a,b]$ (1)
For many neural networks, activation functions, especially the ReLU function, are the main part requiring approximation and naturally is the main cause of approximation errors. ReLu is typically approximated using the formula $\operatorname{ReLU}(z)=\frac{z+z\operatorname{sgn}(z)}{2}$. If the error for $\operatorname{Appsgn}(z)$ is $\epsilon_{sgn}$, then the error for ReLU approximation $\operatorname{AppReLU}(z)=\frac{z+z\operatorname{Appsgn}(z)}{2}$ is:
$\epsilon_{relu}:=\frac{z+z\operatorname{Appsgn}(z)}{2}-\operatorname{ReLU}(z)=\frac{\epsilon_{sgn}\cdot z}{2}$ (2)
If $\operatorname{Appsgn}(z)$ is $\beta$-close, the approximation of $\operatorname{ReLu}(z)$ will have error bound $2^{-(\beta+1)}z$. Equation 2 shows an important fact that the larger the input $z$, the larger the approximation error $\epsilon_{relu}$.
It is also worth noting that $p(z)$ is composed by small degree polynomials: $p=p_{k}\circ p_{k-1}\circ\dots\circ p_{1}$. These small degree polynomials are chosen according to precision. Therefore, PANN with different precision uses different polynomial approximations, leading to different error locations. Therefore, it is hard to improve PANN across different precision simultaneously via fune-tuning.
3 Effects of Approximation Errors on PANN
This section presents our findings regarding approximation errors on neural networks, highlighting its similarities and differences with perturbations in adversarial samples (detailed in Appendix B), and also how neural networks are affected by weight regularization when approximation errors ¡0are involved.
3.1 Perturbations on Different Information
Figure 2: Minimax approximation errors for $\operatorname{Appsgn}(z)$ on $[-1,-\epsilon]$ with precision $2^{-10}$ (left) and $2^{-14}$ (right)
Activation functions are the main components requiring approximation in existing PANNs. Thus, we focus our analyses on the error introduced by ReLU, the most common activation function. As discussed in Section 2.2, approximation of ReLU, $\operatorname{AppReLu}$, is computed by $\frac{z+z\operatorname{Appsgn}(z)}{2}$.The approximation error of $\operatorname{Appsgn}$ function in $[-1,-\epsilon]$ is shown in Figure 2. We note that since $\operatorname{Appsgn}(z)$ is odd, the same approximation errors will also be observed on $[1,\epsilon]$. Assume the error for approximating $\operatorname{sgn}(z)$ is $\epsilon_{sgn}\in[-2^{-\beta},2^{\beta}]$ , we have:
$\epsilon_{relu}:=\operatorname{AppReLU}(z)-\operatorname{ReLU}(z)=\frac{\epsilon_{sgn}\cdot z}{2}$ (3)
We discuss a subtle issue regarding approximation error in PANN involving ReLU. In the backbone model $\operatorname{ReLU(z)}=0$ when $z<0$. In other words, perturbations added to input with $z<0$ will not affect the next layer, whereas approximation error introduced with $z<0$ will be propagated to the next layer. In more detail, we denote the inputs to $\operatorname{ReLU}$ and $\operatorname{AppReLU}$ in a neural network as $z=\boldsymbol{W}_{i}x$. We consider the following two cases: $\boldsymbol{W}_{i}x>0$ and $\boldsymbol{W}_{i}x<0$.
When $\boldsymbol{W}_{i}x>0$, $\operatorname{ReLU}(\boldsymbol{W}_{i}x)=\boldsymbol{W}_{i}x$ in the backbone model, the information of $x$ will be passed to the next layer and may contribute to the final results. We call this “information contributing to the outputs”. For this information, $\epsilon_{relu}=\frac{\epsilon_{sgn}\cdot\boldsymbol{W}_{i}x}{2}$. Since neural networks are nearly linear within a very small interval, we have:
$\operatorname{AppReLU}(\boldsymbol{W}_{i}x)$ $=(1+\frac{\epsilon_{sgn}}{2})\boldsymbol{W}_{i}x$ (4)
$=\operatorname{ReLU}(\boldsymbol{W}_{i}(x+\frac{\epsilon_{sgn}}{2}x))$
If these approximation errors lead to a mistake in the output, their effect is the same as an adversarial sample with perturbation $\frac{\epsilon_{sgn}}{2}x$ on the inputs $x$.
When $\boldsymbol{W}_{i}x<0$, $\operatorname{ReLU}(\boldsymbol{W}_{i}x)=0$ always holds. The information will not be passed to the next layer and contribute to outputs, which we mention as “irrelevant information in the input background”. For this information, we have:
$\operatorname{AppReLU}(\boldsymbol{W}_{i}x)=\frac{\epsilon_{sgn}}{2}\boldsymbol{W}_{i}x=\epsilon_{relu}$ (5)
Different from the previous case, if these errors lead to a wrong output, they can hardly be related to adversarial samples. Since perturbations in adversarial samples are very small and $z$ may be far from 00, it is difficult to make such a large change in $z$ by a finite input perturbation.
In short, the effects of approximation errors in the former case is similar to an evasion attack. In contrast, the second case is unique to the approximation errors in PANNs. We note one more difference, approximation errors differ from adversarial samples as they do not follow the directions of gradients, meaning that while some errors are harmful, some can be beneficial. Nonetheless, approximation error in the second case is most likely harmful to the PANNs because as long as $\boldsymbol{W}_{i}x\neq 0$, errors will lead to unknown behavior in neural networks.
3.2 Observation through Input Perturbations
As discussed, approximation errors affect information contributing to the outputs and irrelevant information in the input background. The effects of approximation errors in the former case are similar to an evasion attack and provide an explanation on how they affect inference accuracy in PANNs. However, it is unclear if approximation errors in the latter case also affect PANNs’ accuracy.
To further investigate the issue, we make use of the fact that $\epsilon_{relu}$ contains the product $\frac{\epsilon_{sgn}\cdot\boldsymbol{W}_{i}x}{2}$. More concretely, we can adjust the magnitude of $\epsilon_{relu}$ by controlling the magnitude of $\boldsymbol{W}_{i}x$. We can seperate the effect of the two cases by adding perturbations to $x$ with $\boldsymbol{W}_{i}x>0$ or $\boldsymbol{W}_{i}x<0$.
For information contributing to the outputs, we can add perturbations to the part of $x$ with $\boldsymbol{W}_{i}x>0$, which is usually the main part of inputs (e.g., the objects in images). If there is an adversarial sample $x+\delta_{app}$ for PANN, we get:
$\operatorname{AppReLU}(\boldsymbol{W}_{i}(x+\delta_{app}))$ (6)
$=(1+\frac{\epsilon_{sgn}}{2})\boldsymbol{W}_{i}(x+\delta_{app})$
$=\operatorname{ReLU}(\boldsymbol{W}_{i}(x+\delta_{app}+\frac{\epsilon_{sgn}}{2}(x+\delta_{app}))$
It means if a valid input perturbation $\delta_{app}$ for PANN exists, we may find a $\delta_{bb}=\delta_{app}+\frac{\epsilon_{sgn}}{2}(x+\delta_{app})$ for the backbone model. Since $\delta_{app}$ has driven $x$ to its class margin and $\frac{\epsilon_{sgn}}{2}$ is very small, $\delta_{bb}$ is easy to find by slightly enlarging $\delta_{app}$.
For irrelevant information in the input background, we can add perturbations to the part of $x$ with $\boldsymbol{W}_{i}x<0$. This part is generally the background of inputs. Different from the previous case, if we find a valid $x_{adv}=x+\delta_{app}$ for PANN here, we have:
$\operatorname{AppReLU}(\boldsymbol{W}_{i}(x+\delta_{app}))=\boldsymbol{W}_{i}\frac{\epsilon_{sgn}}{2}(x+\delta_{app})$ (7)
This error is unreachable by $\operatorname{ReLU}(\boldsymbol{W}_{i}(x+\delta_{bb}))$ for a small $\delta_{bb}$, because $\operatorname{ReLU}$ always outputs 0 for negative inputs. If $\operatorname{AppReLU}(\boldsymbol{W}_{i}(x+\delta_{app}))>0$, we need $\operatorname{ReLU}(\boldsymbol{W}_{i}(x+\delta_{bb}))>0$ then $\boldsymbol{W}_{i}\delta_{bb}>-\boldsymbol{W}_{i}x$, which may require a large $|\delta_{bb}|$; Moreover, if $\operatorname{AppReLU}(\boldsymbol{W}_{i}(x+\delta_{app}))<0$, it’s impossible to find a $\delta_{bb}$ such that $\operatorname{ReLU}(\boldsymbol{W}_{i}(x+\delta_{bb}))<0$. If $\delta_{app}$ can easily mislead PANN, we can identify that approximation errors on irrelevant information in the input background can affect PANN significantly. Meanwhile, if we can hardly find $\delta_{app}$ works on the backbone, we can conclude that this information is irrelevant to the task neural networks are trained for, and that crafting perturbations to this part during training has little damage to the accuracy of the backbone model.
3.3 Effects of weight regularization on PANN
Figure 3: Backbone models (bb) trained with different weight decay (wd) have similar accuracy (dot line). However, large weight decay can reduce their accuracy on PANN as the epoch increases. (ResNet-20, CIFAR-10, precision $2^{-8}$)
Weight regularization is crucial in model training. It can improve model generalization and accuracy Zhang et al. (2018a). However, we observed that it is very harmful to PANN. Even a small weight regularization can significantly reduce the “sturdiness” of the neural networks. An example of PANN on models trained with different weight decay is presented in Figure 3. Their backbone (bb) models have similar accuracy, but the performances of their corresponding PANNs vary greatly. In the early training stages, all these models have a solid resistance to approximation errors, but this resistance is destroyed by weight regularization as the epochs increase. Larger weight decay can make this reduction quicker. At epoch 200, the model with weight decay 0 only reduced the accuracy by about 10%, but weight decay 0.001 reduced the accuracy by over 80%. This reduction is observed in many model structures and has become an important factor affecting the accuracy of PANN (shown in Table 2 and Table 3).
We attribute this to the fact that weight regularization can make the landscape of neural nets more complicated Ziyin et al. (2022). This complexity may not be problematic for normal models, but it can be detrimental to PANN when combined with approximation errors. Though weight regularization can improve generalization within input distribution, approximation errors, especially the errors on irrelevant information in the input background, can be far away from this scope, rendering them useless for PANN.
4 Improving low-precision PANN’s accuracy
This section introduces our methods for improving the “sturdiness” of neural networks and thus improving low-precision PANN’s accuracy.
4.1 AT-like Solutions for Improving PANN
We propose a solution to enhance “sturdiness” based on the findings about the relationship between approximation error and adversarial robustness discussed in Section 3.1. This inspire us to adapt adversarial training to deal with approximation errors.
Training Methods: For a neural network with layered architectures, the input layer is numbered 00, the output layer is numbered $L$, and the hidden layers $h_{l}$ are numbered $1$, $\dots,L-1$. We denote $z_{l}$ as the input at layer $l$ before the activation function and $z_{L}$ is the output of the neural network, $x_{l}$ as the vector after application of the activation function (to $z_{l-1}$) and $x_{0}=x$ is the neural network input. For normal models, $z_{l}=h_{l}(x_{l})$. For our method, we use $\mathbb{F}^{\prime}(x)$ to represent the $\mathbb{F}(x)$ replacing all $z_{l}$ with $z^{\prime}_{l}=h_{l}(x_{l}+\epsilon_{in,l})$, use $\beta^{\prime}$ to represent the amplification due to the multiplication of errors with intermediate values. We are looking for a neural network that satisfies the following:
$\inf_{\theta}\mathbb{E}_{\mathbb{P}}[\mathcal{L}(\mathbb{F}^{\prime}(x),y)],\quad\forall||\epsilon_{in,l}||<2^{-\beta+\beta^{\prime}}$ (8)
When trying to design this AT-like method, there are two considerations. First, adversarial samples only affect information contributing to the outputs, but we analyze and prove that approximation errors can also affect irrelevant information in the input backgrounds. Second, adversarial samples only add perturbations to inputs, while approximation errors are generated in almost every layer. These differences result in a much wider range of perturbations in PANN, with a much larger amplitude (after accumulated and amplified). Therefore, we should not introduce perturbations during training at every point susceptible to approximation errors like AT does because the accuracy of the backbone model will be reduced. This reduction will limit the upper bound of PANN accuracy and affect the use of the backbone model in other areas, thereby increasing the overall costs.
To address this issue, we consider restricting perturbations only to specific positions. According to Equation 3, approximation errors are multiplied by function inputs, which means that errors on activation function inputs with large absolute values are more detrimental. Therefore, we can add perturbations only to the proportion of activation function inputs with the largest magnitude. Furthermore, intra-model perturbations consist of two parts with different characters. One part affects information contributing to the outputs and can be reached by small perturbations on the backbone model input (usually the positive inputs to ReLU), denoted by $e_{l}(x+\delta_{bb})$. The other part affects irrelevant information in the input background and is unreachable by $\delta_{bb}$ (usually the negative inputs to ReLU), donated by $\epsilon_{irr,l}$. So we have $\epsilon_{int,l}=e_{l}(x+\delta_{bb})+\epsilon_{irr,l}$. The first part does not follow the gradient direction and is not effective enough. In contrast, the second part is always harmful (explained in Section 3.1). This means that adding perturbations to negative values with large magnitudes is the most effective and has minimal effect on the backbone model’s performance.
With the reasons above, we design a noise generator for negative values with large magnitude (NGNV) in activation function inputs during the training phase. For the smallest $r$ (eg: 0.7) percent negative values in the ReLU inputs, we generate Gaussian noise $\epsilon_{gau}$ to simulate $\epsilon_{sgn}$ and multiply with these values. The products are then multiplied by a parameter $\lambda$ (eg: 0.05) and added to ReLU inputs. In other words, for a ReLU input $z$, we generate intra-model perturbations $\delta_{int}$ as:
$\delta_{int}=\lambda\epsilon_{gau}z,\quad\epsilon_{gau}\leftarrow\mathcal{N}(0,1)$ (9)
Considering that the approximation error $\epsilon_{relu}$ has boundaries. We can extract signs from $\epsilon_{g}$ and simulate the worst case by setting a fixed $\lambda$. The process is:
$\delta_{int}=\lambda\operatorname{sgn}(\epsilon_{gau})z,\quad\epsilon_{gau}\leftarrow\mathcal{N}(0,1)$ (10)
4.2 Reduced Use of Weight Regularization
As discussed in Section 3.3, weight regularization reduces PANNs’ accuracy. One possible solution is to use minimal weight regularization combined with early stopping. However, it will lead to poor generalization, which can reduce the accuracy of backbone models, thereby limiting PANN’s accuracy by reducing its upper bound.
Therefore, we choose Mixup as a remedy method. Mixup Zhang et al. (2018b) introduces globally linear behavior in-between data manifolds. It can reduce unexpected behavior of neural networks and enhance generalization and robustness by expanding the input set. Our experiments show that Mixup can somewhat compensate for the reduced accuracy due to poor regularization. The test results based on minimal regularization and Mixup are presented in Section 5.2, which shows good performance on some model structures and precision. However, it is important to note that this approach is only effective when regularization is not crucial. For deeper models, the problem of low backbone accuracy under poor regularization is still challenging. Besides, Mixup may decrease PANN’s accuracy for certain models, such as Shufflenetv2. Thus, we regard use of low weight regularization and Mixup as a trade-off.
5 Experiments
We conduct experiments to demonstrate effects of the perturbations on irrelevant information in the input background as well as the effectiveness of our solutions to improve sturdiness and the accuracy of low-precision PANN. The main results are shown in this section. Additional experiment results can be found in Appendix A.
5.1 Effects of Approximation Errors illustrated by PGD attacks
(a) Success rate of PGD attacks on PANN
(b) Percentage of $\delta$ only affects PANN
Figure 4: (a) PGD attacks on both regions can have high success rates on PANN. (b)PGD attacks on large-gradient regions have high success rates in backbone models, but attacks on small-gradient regions can hardly affect the backbone model.
Section 3.2 mentioned that approximation errors lead to intra-model perturbations on irrelevant information in the input background and are extremely harmful to PANN. To prove that, we create adversarial samples with perturbation added to irrelevant information in the input background and information contributing to the outputs separately. “Irrelevant” means the information is filtered and has little effect on model outputs and corresponds to image background in images. If we create adversarial samples with perturbation added to input backgrounds, the perturbations will be transferred throughout PANN by approximation errors and enlarge the error by multiplication. This will severely impact PANN but have little effect on the backbone models. However, if we attack objects in images (corresponding to information contributing to the outputs), both PANN and backbone will be affected.
In our experiment, we create adversarial samples with perturbation added to objects and backgrounds separately by preserving part of input perturbations $\delta$ according to magnitude of gradients in the backbone model: Attacks on large-gradient regions preserve perturbations on pixels with the top 20% gradient magnitude (usually the object); Attacks against small-gradient regions preserve pixels with the minimum 30% gradient magnitude (usually the image background). For each $\alpha$, we randomly select 5000 images.
Figure 4 shows that attacks on backgrounds and objects all have a high success rate on PANN. However, only attacks on objects can easily affect backbone models. When the step length is very small, attacks on both large-gradient and small-gradient regions will first only mislead PANN. Immediately after the step length increases, attacks on large-gradient regions also affect the backbone model. In contrast, attacks on small-gradient regions only have 32% success rate on backbone models even though the step length is 40 times the beginning. This result is a good proof of the assumption in Section 3.2. It demonstrates that PANN suffers from approximation errors in the irrelevant information in the input background, opening the possibility of dealing with approximation errors by leveraging this information.
5.2 Accuracy Evaluation
Table 2: Accuracy of PANN on ResNet-20, Shufflenetv2, DLA-34, Mobilenetv2 trained with weight decay (wd) 0, 1e-4 and 5e-4. We give Top-1 accuracy of the backbone (bb) and PANN with precision $2^{-\beta}$ ($2^{-8}$ and $2^{-9}$). Training methods include Vanilla, Mixup, and robustness-based solution NGNV. The results with the highest accuracy for each model and approximation precision are bolded.
Vanilla Mixup NGNV Mixup+NGNV
$\beta$ wd 0 1e-4 5e-4 0 1e-4 5e-4 0 1e-4 5e-4 0 1e-4 5e-4
[…]
bb 90.39 91.57 92.14 91.30 91.60 92.17 91.10 91.70 92.49 91.49 91.59 92.16
[…]
bb 88.54 90.22 88.60 89.08 90.19 87.62 88.66 90.69 89.60 89.10 90.54 88.32
[…]
bb 93.43 94.38 95.10 94.96 95.12 95.41 93.85 94.92 95.21 95.09 95.62 95.58
[…]
bb 92.69 93.49 91.45 93.60 93.74 90.50 92.97 93.58 91.84 93.68 94.03 91.58
Table 3: Accuracy of PANN with precision $2^{-\beta}$ and the backbone (bb) on CIFAR-100 (ResNet-32) and TinyImagenet (ResNet-18).
Vanilla Mixup Mixup+NGNV
$\beta$ wd 0 1e-4 0 1e-4 0 1e-4
[…]
bb 65.08 68.88 67.92 70.96 68.85 70.32
[…]
bb 56.62 59.62 58.78 62.01 59.42 61.30
Evaluation Setup: We conduct tests on PANN with ResNet He et al. (2015), DLA Yu et al. (2018), MobilenetV2 Sandler et al. (2019), and ShufflenetV2 Ma et al. (2018) models for classification tasks on CIFAR-10. We also test ResNet on CIFAR-100 and Tiny Imagenet dataset Krizhevsky et al. (2009); Le & Yang (2015). Training on CIFAR dataset takes 200 epochs, and on Tiny Imagenet takes 150 epochs. All learning rates begin with 0.1, and were multiplied by 0.1 on epochs 100 and 150 on CIFAR dataset, multiplied by 0.1 on epochs 50 and 100 on Tiny Imagenet. The momentum is 0.9. We test weight decay on 0, 0.0001, and 0.0005. Note that the ResNet-18 models removed the Maxpool layer to suit small images. $\lambda$ for Mixup are selected from distribution $\operatorname{Beta}(0.5,0.5)$. The approximation intervals are $[-50,50]$ for CIFAR and $[-100,100]$ for Tiny Imagenet, except the case values exceed the approximation interval. The average accuracy of the backbone models and PANN with precision $2^{-8}$ and $2^{-9}$ on CIFAR-10 dataset are presented in Table 2. The experiments on CIFAR-100 (C100) and Tiny Imagenet are shown in Table 3. The highest accuracy for each model and approximation precision are highlighted in bold.
Effects of Weight Regularization: From Table 2 and Table 3, we observe that weight regularization can severely destroy the model’s resistance to approximation errors in various models and precisions. For example, PANN on models trained with weight decay 0 outperforms those with weight decay 5e-4 in nearly all vanilla models, suggesting minimal weight decay. Note that this discipline doesn’t always hold. Poor regularization caused by no regularization can limit backbone models’ accuracy, thus reducing PANN’s accuracy by limiting the upper bound. For example, ResNet-20 models with weight decay 1e-4 can outperform those with zero weight decay in precision $2^{-9}$.
Evaluation of Minimal Regularization with Mixup: Table 2 and Table 3 shows that Mixup can effectively compensate for the lack of generalization and reduced backbone accuracy caused by poor regularization. Therefore, models can benefit from strong resistance to approximation errors which come with minimal weight regularization. This method can improve PANN performance in many cases. For example, Mixup can outperform vanilla models on precision $2^{-9}$ with zero weight decay in ResNet-20 and DLA-34. However, this advantage may not always hold. In some cases, Mixup might be harmful to PANN (eg: DLA-34, weight decay 0, precision $2^{-8}$).
Evaluation of AT-like Solution: Table 2 and Table 3 prove that NGNV (or NGNV with Mixup), can achieve the best accuracy in all precisions (bolded). It can also improve PANN accuracy significantly in almost all cases, even if the weight decay is large. For instance, on Shufflenetv2 and Mobilenetv2, our models can improve PANN accuracy by up to 60% on weight decay 0 and precision $2^{-8}$, by over 40% on weight decay 1e-4 and precision $2^{-9}$.
In a word, by combining these methods, models can exhibit strong “sturdiness” to approximation errors. We can significantly improve low-precision PANN’s accuracy and thus reduce the time cost. Table 1 illustrates some time costs for PANN with different precision on different models. Previous schemes require at least $2^{-12}$ precision to achieve satisfactory accuracy, but our method only needs $2^{-9}$ (Kindly refer to the appendix for the selection of parameters). The time required for PANN with our models to achieve the state-of-the-art Lee et al. (2021a; 2022b) accuracy is given in Figure 1. Our solution can reduce the PANN time cost by 40% to 60% compared to models provided by these schemes. This improvement makes it easier to design cryptographic schemes for various scenarios.
6 Conclusions
We uncovered crucial insights into the nature of “sturdiness”, a notion we introduced as a network’s resistance to approximation errors. We identified the intricate relationship between “sturdiness” and adversarial robustness and introduce an AT-like method tailored to enhance sturdiness. Additionally, we discovered that weight regularization negatively affect “sturdiness”, and explored alternative generalization strategies.
We believe this work represents an important step for advancements in PANNs. Our findings will encourage continued research, which we believe will lead to more efficient and effective PPML.
We conclude with a few open problems that warrant deeper exploration, namely, establishing a precise mathematical relationship between model accuracy and approximation error, understanding the implications of weight regularization, and design of solutions applicable at different precision levels.
Appendix A Additional Experiment Results
Table 4: Accuracy of PANN on CIFAR-10 dataset (ResNet-20, Shufflenetv2, DLA-34, Mobilenetv2) trained with weight decay (wd) 0, 1e-4 and 5e-4. We give Top-1 accuracy of the backbone (bb) and PANN with precision $2^{-\beta}$. Training methods include Vanilla, Mixup, and robustness-based solution NGNV. The results with the highest accuracy for each model and approximation precision are bolded.
Vanilla Mixup NGNV Mixup+NGNV
$\beta$ wd 0 1e-4 5e-4 0 1e-4 5e-4 0 1e-4 5e-4 0 1e-4 5e-4
[…]
bb 90.39 91.56 92.14 91.30 91.60 92.17 91.10 91.69 92.49 91.49 91.59 92.16
[…]
bb 88.54 90.22 88.60 89.08 90.19 87.62 88.66 90.69 89.60 89.10 90.54 88.32
[…]
bb 93.43 94.38 95.10 94.96 95.12 95.41 93.85 94.92 95.21 95.09 95.62 95.58
[…]
bb 92.69 93.49 91.45 93.60 93.74 90.50 92.97 93.58 91.84 93.68 94.03 91.58
This section gives more results of the accuracy of PANN on models trained with different methods and weight decays. The experiment results on CIFAR-10 dataset are shown in Table 4, on CIFAR-100 (C100) and TinyImagenet is shown in Table 6. Note that PANN’s accuracy is very close to the backbone models for high precision like $2^{-11}$ and $2^{-12}$, which reduce the advantage of small weight regularization.
Table 5: Accuracy and time cost of PANN with precision $2^{-\beta}$ and the backbone (bb) on CIFAR-100 (ResNet-32) and TinyImagenet (ResNet-18). The results with the highest accuracy for each model and approximation precision are bolded.
Vanilla Mixup Mixup+NGNV Time Cost
$\beta$ wd 0 1e-4 5e-4 0 1e-4 5e-4 0 1e-4 5e-4
[…]
bb 65.08 68.88 70.30 67.92 70.96 70.39 68.85 70.32 69.31 2.5s
[…]
bb 56.62 59.62 62.17 58.78 62.01 63.47 59.42 61.30 61.77 24.7s
We also give parameters $r$ and $\lambda$ for our NGNV in Table 6 to help the reader reproduce our results.
Table 6: Best parameters ($r$, $\lambda$) for NGNV on different weight decay and methods
NGNV Mixup+NGNV
wd 0 1e-4 5e-4 0 1e-4 5e-4
CIFAR-10 ResNet-20 (0.3,0.3) (0.1,0.3) (0.3,0.2) (0.5,0.1) (0.5,0.1) (0.5,0.05)
Shufflenetv2 (0.5,0.05) (0.1,0.3) (0.6,0.1) (0.5,0.1) (0.3,0.05) (0.5,0.1)
DLA-34 (0.1,0.3) (0.1,0.3) (0.5,0.05) (0.3,0.05) (0.3,0.3) (0.1,0.3)
Mobilenetv2 (0.5,0.1) (0.5,0.05) (0.5,0.05) (0.3,0.05) (0.5,0.1) (0.3,0.1)
CIFAR-100 ResNet-32 (0.1,0.3) (0.3,0.3) (0.1,0.3)
Tiny-Imagenet ResNet-18 (0.3,0.3) (0.1,0.3) (0.3,0.3)
The accuracy and time cost for models released by us and the state-of-the-art schemes on different precisions are presented in Table 7. (Time cost for ResNet-20 has been given in Table 1.)
Table 7: Comparation of Models Provided by us and Models offered by the State-of-the-Art Scheme Lee et al. (2022b) (Classification Accuracy on CIFAR-10)
$\beta$ 8 9 10 11 12 bb
State-of-the-Art 31.79 87.57 90.52 91.22 91.43 91.51
ResNet-20 Ours 90.49 91.74 91.88 92.03 92.05 92.02
State-of-the-Art 35.71 88.32 91.85 92.25 92.39 92.48
Ours 90.55 91.92 92.42 92.42 92.42 92.42
ResNet-32 Time cost 45.49s 53.54s 60.81s 97.88s 105.33s 2.87s
State-of-the-Art 11.37 60.75 91.09 92.55 92.75 92.75
Ours 72.27 90.30 92.85 93.21 93.27 93.34
ResNet-44 Time cost 56.68s 68.57s 86.43s 133.87 149.66s 2.90s
State-of-the-Art 11.05 42.54 90.12 92.75 93.12 93.26
Ours 60.54 89.50 93.30 93.63 93.77 93.81
ResNet-56 Time cost 79.32s 93.98s 110.73s 173.72s 189.85s 3.42s
State-of-the-Art 10.73 17.72 71.27 92.01 93.22 93.49
Ours 49.90 84.65 93.68 94.04 94.16 94.16
ResNet110 Time cost 148.25s 181.37s 215.87s 348.38s 362.27s 4.01s
Appendix B Introduction to Adversarial Samples, Adversarial Training and Mixup
B.1 Adversarial Samples and Adversarial Training
Adversarial sample are used in evaison attacks to deceive neural networks. They can be generated by taking small steps in the direction of the loss gradient Goodfellow et al. (2014). A commonly used methods to generate adversarial samples is Projected Gradient Descent Attack (PGD) Kurakin et al. (2016). PGD searches for adversarial samples interactively in a small area around the input by adding perturbations following loss gradients:
$x^{adv}_{t+1}=x^{adv}_{t}+\alpha\cdot sign(\nabla_{x}\mathcal{L}(\mathbb{F}(x^{adv}_{t}),y))$ (11)
Adversarial training (AT) was first proposed by PGD-AT to increase the model’s robustness to adversarial samples Madry et al. (2018). Adversarial robustness prevents neural networks from being fooled, and can be achieved by ensuring that all samples near the sample $x$ have the same classification with $x$: $\mathbb{F}(x)=\mathbb{F}(x+\delta)$ for $\forall|\delta|<\epsilon$. AT introduces adversarial samples to the training set, seeking the worst case in scope $\ B_{\epsilon}(x)=\{x^{\prime}:c_{\mathcal{X}}(x,x^{\prime})\leq\epsilon)\}$:
$\inf_{\theta}\mathbb{E}_{\mathbb{P}}[\sup_{x^{\prime}\in B_{\epsilon}(x)}\mathcal{L}(\mathbb{F}(x^{\prime}),y)]$ (12)
B.2 Mixup
Mixup is the method that introduces interpolated samples to the training set Zhang et al. (2018b), which was proved to help model robustness and generalization Zhang & Deng (2021). For parameter $\lambda$ and samples $(x_{i},y_{i})$ and $(x_{j},y_{j})$ randomly selected from the training set, the interpolated sample is:
$\widetilde{x}=\lambda x_{i}+(1-\lambda)x_{j},\quad\widetilde{y}=\lambda y_{i}+(1-\lambda)y_{j}$ (13)
Appendix C Adversarial Samples against PANN only
We present a method to generate adversarial samples against PANN only to help readers better observe their differences. An intuitive way to find adversarial samples only for PANN is to look for pixels where gradients differ most. However, information contributing to the outputs also has gradient differences. As discussed in Section 3, perturbations on this information can affect both models. So, we set a small value $\epsilon_{lim}$ to limit the gradient changes on the backbone model. Besides, to make the phenomenon more apparent, we set $\epsilon_{atk}$ to preserve only large gradient differences.
Figure 5: The benign samples and perturbations $\delta$ only against PANN. $\delta$ can concentrate on image backgrounds (ResNet-18, Imagenet, Precision $2^{-14}$)
Algorithm 1 Evasion Attacks against PANN Only
Input: Clean Input $x$;
Parameter: Backbone model $\mathbb{F}$; PANN $\widetilde{\mathbb{F}}$; Attack Step Length $\alpha$; Limitation for Perturbations $\epsilon$
Output: Perturbation $\delta$
1:  Initialize $\delta_{0}$ as all zero;
2:  while $\mathbb{F}(x+\delta)\neq y$ or $\widetilde{\mathbb{F}}(x+\delta)=y$ do
3:     if $\widetilde{\mathbb{F}}(x+\delta)=y$ then
4:        $\delta=\delta+\alpha\nabla_{\delta}\mathcal{L}_{\widetilde{\mathbb{F}}}(x+\delta)$
5:        $\delta=Clip_{[-\epsilon,\epsilon]}(Random\ Search(x+\delta))$
6:        $\delta=\delta\odot Mask(\nabla_{\delta}(\mathcal{L}_{\widetilde{\mathbb{F}}}(x+\delta)-\nabla_{\delta}\mathcal{L}_{\mathbb{F}}(x+\delta))\geq\epsilon_{atk})$
7:        $\delta=\delta\odot Mask(|\nabla_{\delta}\mathcal{L}_{\mathbb{F}}(x+\delta)-\nabla_{\delta}\mathcal{L}_{\mathbb{F}}(x)|\leq\epsilon_{lim})$
8:     end if
9:     if $\mathbb{F}(x+\delta)\neq y$ then
10:        $Backtrack(\delta)$
11:     end if
12:  end while
13:  Return $\delta$
Additionally, two problems were faced when using PGD-based attacks: (1)Firm-step search can cross the adversarial samples. As shown in Figure 2, approximation errors fluctuate when the input changes, so valid perturbations are in discrete intervals. It’s difficult to estimate where these intervals are. Using particularly smaller steps can avoid this, but is inefficient. Therefore, we apply random searches after each fixed step. (2) Even if limits are set, the perturbations may still affect the backbone model. Once the backbone model gives a wrong output, the PGD interaction will fall into this area and can hardly escape. To avoid this, we apply backtracking once the backbone model gives the wrong results. The whole process is represented in Algorithm 1.
We attack PANN with precision $2^{-14}$ and approximation interval $[-100,100]$. Some results on the Pytorch Resnet-18 pre-trained model are shown in Figure 5. It can be seen that $\delta$ can concentrate on both backgrounds and objects. Especially when $\alpha$ is not too small, $\delta$ can locate mainly on backgrounds for many samples, which provides evidence for the differences in irrelevant information in the input background between PANN and backbone models.